# Parabolic induction in the homotopy category of pro-$p$ Iwahori-Hecke modules

### v2

@@ 4.1. Constructing model structures from a recollement
(4.1) ${{\mathcal{S}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{i}$${{\operatorname{Ho}}({\mathcal{A}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\mathsf{\pi_{\mathcal{S}}}}$$\scriptstyle{\mathbf{L}{\mathsf{L}}}$$\scriptstyle{\mathbf{R}{\mathsf{R}}}$${{\operatorname{Ho}}({\mathcal{A}})/{\mathcal{S}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\ell_{\mathcal{S}}}$$\scriptstyle{r_{\mathcal{S}}}$
where the functors $\mathbf{L}{\mathsf{L}}$ and $\mathbf{R}{\mathsf{R}}$ are not a priori assumed to be the total derived functor of some Quillen functors.
We recall that the class of trivially fibrant objects in an injective abelian model category equals the class $\mathrm{Inj}({\mathcal{A}})$ of injective objects, and dually the class of trivially cofibrant objects in a projective abelian model category equals the class $\mathrm{Proj}({\mathcal{A}})$ projective objects (c.f. [7, Corollary 1.1.9]).
@@
Let ${\mathcal{A}}$ be a Grothendieck abelian category equipped with a cofibrantly generated, abelian model structure which is either injective or projective. Suppose that ${\mathcal{S}}$ is a thick subcategory of ${\operatorname{Ho}}({\mathcal{A}})$ satisfying 4.1. Let ${\mathcal{W}}_{\mathcal{S}}$, resp. ${\mathcal{F}}_{\mathcal{S}}$, resp. ${\mathcal{C}}_{\mathcal{S}}$, denote the class of objects of ${\mathcal{A}}$ whose image in ${\operatorname{Ho}}({\mathcal{A}})$ lies in ${\mathcal{S}}$, resp. ${\operatorname{Ker}}(\mathbf{R}{\mathsf{R}})$, resp. ${\operatorname{Ker}}(\mathbf{L}{\mathsf{L}})$.
@@
Since distinguished triangles arise from short exact sequences, the result follows immediately from the fact that ${\mathcal{S}}$, ${\operatorname{Ker}}(\mathbf{R}{\mathsf{R}})$, and ${\operatorname{Ker}}(\mathbf{L}{\mathsf{L}})$ are thick subcategories of the triangulated category ${\operatorname{Ho}}({\mathcal{A}})$. ∎
@@
Suppose that we’re in 4.1. When ${\mathcal{A}}$ is injective then ${\mathcal{W}}_{\mathcal{S}}\cap{\mathcal{F}}_{\mathcal{S}}={\mathcal{W}}_{\mathcal{A}}$, so that ${\mathcal{W}}_{\mathcal{S}}\cap{\mathcal{F}}_{\mathcal{A}}\cap{\mathcal{F}}_{\mathcal{S}}={\mathcal{W}}_{\mathcal{A}}\cap{\mathcal{F}}_{\mathcal{A}}=\mathrm{Inj}({\mathcal{A}})$. Dually, when ${\mathcal{A}}$ is projective then ${\mathcal{W}}_{\mathcal{S}}\cap{\mathcal{C}}_{\mathcal{S}}={\mathcal{W}}_{\mathcal{A}}$, so that ${\mathcal{W}}_{\mathcal{S}}\cap{\mathcal{C}}_{\mathcal{A}}\cap{\mathcal{C}}_{\mathcal{S}}={\mathcal{W}}_{\mathcal{A}}\cap{\mathcal{C}}_{\mathcal{A}}=\mathrm{Proj}({\mathcal{A}})$.
@@
We only prove the injective case, the other one being dual. Note that ${\mathcal{W}}_{\mathcal{A}}\cap{\mathcal{F}}_{\mathcal{A}}=\mathrm{Inj}({\mathcal{A}})$ is immediate from the injectivity of the model structure on ${\mathcal{A}}$, so that the first equality immediately implies the second. For the first equality, we have ${\mathcal{W}}_{\mathcal{A}}\subseteq{\mathcal{W}}_{\mathcal{S}}\cap{\mathcal{F}}_{\mathcal{S}}$ by definition of the right hand side. Conversely, assume that $X\in{\mathcal{W}}_{\mathcal{S}}\cap{\mathcal{F}}_{\mathcal{S}}$. Then $X\cong i(Y)$ in ${\operatorname{Ho}}({\mathcal{A}})$, for some $Y\in{\mathcal{S}}$, and $X\in{\operatorname{Ker}}(\mathbf{R}{\mathsf{R}})$. But then we obtain from 2.4 that $Y=0$ in ${\operatorname{Ho}}({\mathcal{A}})$, and hence that $X\in{\mathcal{W}}_{\mathcal{A}}$ as required. ∎
Let ${\mathcal{A}}$ be a Grothendieck abelian category. Recall that a cotorsion pair $({\mathcal{D}},{\mathcal{E}})$ on ${\mathcal{A}}$ is said to be cogenerated by a set $\mathcal{G}\subseteq{\mathcal{D}}$ if ${\mathcal{E}}=\mathcal{G}^{\perp}$. Furthermore, by definition ${\mathcal{A}}$ has a generating set $\mathscr{S}$ (or equivalently a generator), which is a set of objects of ${\mathcal{A}}$ such that every object of ${\mathcal{A}}$ is a quotient of a direct sum of objects in $\mathscr{S}$. A cotorsion pair on ${\mathcal{A}}$ is called small if it is cogenerated by a set and if ${\mathcal{D}}$ contains a generating set for ${\mathcal{A}}$. We note that small cotorsion pairs are automatically functorially complete and that an abelian model structure $({\mathcal{C}},{\mathcal{F}},{\mathcal{W}})$ on ${\mathcal{A}}$ is cofibrantly generated if and only if both $({\mathcal{C}},{\mathcal{F}}\cap{\mathcal{W}})$ and $({\mathcal{C}}\cap{\mathcal{W}},{\mathcal{F}})$ are small, c.f. [15, Corollary 6.6] and [7, Proposition 1.2.7].
We will also need a notion of generating set for a triangulated category ${\mathcal{T}}$. A set of objects $\mathcal{G}$ of ${\mathcal{T}}$ will be called a set of weak generators if an object $X\in{\mathcal{T}}$ is zero if and only if ${\operatorname{Hom}}_{\mathcal{T}}(G,\Sigma^{n}X)=0$ for all $n\in{\mathbb{Z}}$ and all $G\in\mathcal{G}$. The relevance of this notion for us is the result [14, Theorem 7.3.1], which in particular implies that the homotopy category ${\operatorname{Ho}}({\mathcal{A}})$ of a hereditary, abelian, cofibrantly generated model category ${\mathcal{A}}$ has a set of weak generators given by the cokernels of the generating cofibrations. The proof of loc. cit. requires some input from the homotopy theory of simplicial sets, so we deemed it worthwhile to provide a more elementary proof in the case where ${\mathcal{A}}$ is the module category of a Gorenstein ring. In fact we obtain a stronger result than weak generation.
@@
We first note that 4.5 implies that $({\mathcal{A}},{\mathcal{W}}_{\mathcal{S}}\cap{\mathcal{F}}_{\mathcal{A}}\cap{\mathcal{F}}_{\mathcal{S}})=({\mathcal{A}},\mathrm{Inj}({\mathcal{A}}))$ in case (i) and $({\mathcal{W}}_{\mathcal{S}}\cap{\mathcal{C}}_{\mathcal{A}}\cap{\mathcal{C}}_{\mathcal{S}},{\mathcal{A}})=(\mathrm{Proj}({\mathcal{A}}),{\mathcal{A}})$ in case (ii) are small cotorsion pairs, c.f. [7, Example 1.2.6]. Since ${\mathcal{W}}_{\mathcal{S}}$ is thick by 4.4, we are left to deal with the remaining two pairs.
@@
(i) We need to construct a set $\mathcal{G}$ of objects of ${\mathcal{A}}$ such that ${\mathcal{F}}_{\mathcal{A}}\cap{\mathcal{F}}_{\mathcal{S}}=\mathcal{G}^{\perp}$ and ${\mathcal{W}}_{\mathcal{S}}={{}^{\perp}}(\mathcal{G}^{\perp})$. By assumption on ${\mathcal{A}}$, we have ${\mathcal{F}}_{\mathcal{A}}=\mathcal{G}_{1}^{\perp}$ for some set $\mathcal{G}_{1}$ of objects in ${\mathcal{W}}_{\mathcal{A}}$. By enlarging $\mathcal{G}_{1}$ if necessary, we may assume that it contains a generating set for ${\mathcal{A}}$, since $({\mathcal{W}}_{\mathcal{A}},{\mathcal{F}}_{\mathcal{A}})$ is small. We now let $\mathscr{S}$ be a set of weak generators for ${\operatorname{Ho}}({\mathcal{A}})$ and, for each $S\in\mathscr{S}$ and each $n\in{\mathbb{Z}}$, we choose $X_{n,S}\in{\mathcal{A}}={\mathcal{C}}_{\mathcal{A}}$ such that $X_{n,S}\cong\Sigma^{-n}i\mathbf{L}{\mathsf{L}}(S)$ in ${\operatorname{Ho}}({\mathcal{A}})$ and let $\mathcal{G}_{2}=\{X_{n,S}\mid S\in\mathscr{S}\text{ and }n\in{\mathbb{Z}}\}$. Then we claim that $\mathcal{G}=\mathcal{G}_{1}\cup\mathcal{G}_{2}$ satisfies our requirements. To see this, first note that if $Y\in\mathcal{G}^{\perp}$ then in particular $Y\in\mathcal{G}_{1}^{\perp}={\mathcal{F}}_{\mathcal{A}}$. So in order to obtain ${\mathcal{F}}_{\mathcal{A}}\cap{\mathcal{F}}_{\mathcal{S}}=\mathcal{G}^{\perp}$ it suffices to show that for $Y\in{\mathcal{F}}_{\mathcal{A}}$, we have $Y\in\mathcal{G}_{2}^{\perp}$ if and only if $Y\in{\mathcal{F}}_{\mathcal{S}}$. Indeed, by 2.16 we have
@@
$Y\in\mathcal{G}_{2}^{\perp}$ $\iff{\operatorname{Hom}}_{{\operatorname{Ho}}({\mathcal{A}})}(\Sigma^{-n}i\mathbf{L}{\mathsf{L}}(S),\Sigma Y)=0\text{ for all }n\in{\mathbb{Z}}\text{ and }S\in\mathscr{S}$
$\iff{\operatorname{Hom}}_{{\operatorname{Ho}}({\mathcal{A}})}(S,\Sigma^{n+1}i\mathbf{R}{\mathsf{R}}(Y))=0\text{ for all }n\in{\mathbb{Z}}\text{ and }S\in\mathscr{S}$
$\iff i\mathbf{R}{\mathsf{R}}(Y)=0\iff Y\in{\mathcal{F}}_{\mathcal{S}}$
@@
as required. Next, note that every object of ${\mathcal{F}}_{\mathcal{S}}$ is isomorphic in ${\operatorname{Ho}}({\mathcal{A}})$ to an object of ${\mathcal{F}}_{\mathcal{A}}\cap{\mathcal{F}}_{\mathcal{S}}$. Thus we see by applying 2.16 again that an object $X\in{\mathcal{A}}$ lies in ${}^{\perp}({\mathcal{F}}_{\mathcal{A}}\cap{\mathcal{F}}_{\mathcal{S}})$ if and only if the image of $X$ in ${\operatorname{Ho}}({\mathcal{A}})$ lies in ${}^{\ddagger}{\operatorname{Ker}}(\mathbf{R}{\mathsf{R}})={\mathcal{S}}$.
@@
This shows that $({\mathcal{W}}_{\mathcal{S}},{\mathcal{F}}_{\mathcal{A}}\cap{\mathcal{F}}_{\mathcal{S}})$ is a small cotorsion pair, and thus we have that ${\mathcal{A}}^{i}_{\mathcal{S}}$ defines an injective, cofibrantly generated, abelian model structure as required. By 4.5 we have that the trivial fibrations in ${\mathcal{A}}$ and in ${\mathcal{A}}^{i}_{\mathcal{S}}$ are the same, and thus ${\operatorname{id}}:{\mathcal{A}}^{i}_{\mathcal{S}}\to{\mathcal{A}}$ is right Quillen. Furthermore, since the trivialy fibrant objects in both model structures are the same we have that $\mathbf{R}{\operatorname{id}}:{\operatorname{Ho}}({\mathcal{A}}^{i}_{\mathcal{S}})\to{\operatorname{Ho}}({\mathcal{A}})$ identifies, after restricting to the cofibrant-fibrant objects, with the canonical embedding $({\mathcal{F}}_{\mathcal{S}}\cap{\mathcal{F}}_{\mathcal{A}})/\mathrm{Inj}({\mathcal{A}})\hookrightarrow{\mathcal{F}}_{\mathcal{A}}/\mathrm{Inj}({\mathcal{A}})$ with image ${\operatorname{Ker}}(\mathbf{R}{\mathsf{R}})\simeq{\operatorname{Ho}}({\mathcal{A}})/{\mathcal{S}}$, where the latter equivalence is induced by $r_{\mathcal{S}}$. Thus, after these identifications, the left adjoint of $\mathbf{R}{\operatorname{id}}$ is naturally isomorphic to $\pi_{\mathcal{S}}$.
@@
(ii) As above, we let $\mathscr{S}$ be a set of weak generators for ${\operatorname{Ho}}({\mathcal{A}})$. Then for each $S\in\mathscr{S}$ and each $n\in{\mathbb{Z}}$, we choose $X_{n,S}\in{\mathcal{C}}_{\mathcal{A}}$ such that $X_{n,S}\cong\Sigma^{-n}\ell_{\mathcal{S}}\pi_{\mathcal{S}}(S)$ in ${\operatorname{Ho}}({\mathcal{A}})$ and let $\mathcal{G}=\{X_{n,S}\mid S\in\mathscr{S}\text{ and }n\in{\mathbb{Z}}\}$. Since the model structure on ${\mathcal{A}}$ is projective, there is a projective generator $Z$ for ${\mathcal{A}}$ and we may apply [15, Corollary 6.9] to see that $\mathcal{G}$ cogenerates a small cotorsion pair $({\mathcal{D}},{\mathcal{E}})$, where ${\mathcal{D}}$ is the closure of $\mathcal{G}\cup\{Z\}$ under transfinite compositions and direct summands. As ${\mathcal{C}}_{\mathcal{A}}$ is closed under those (c.f. Lemma 6.2 in loc. cit.), it follows that ${\mathcal{D}}\subseteq{\mathcal{C}}_{\mathcal{A}}$. We first show that ${\mathcal{E}}=\mathcal{G}^{\perp}={\mathcal{W}}_{\mathcal{S}}$. This is similar to the argument in (i). Since every object of $\mathcal{G}$ is cofibrant, we may apply 2.16 to see that
@@
as required, where the last equivalence holds because ${\operatorname{Ker}}(\pi_{\mathcal{S}})={\mathcal{S}}$. Next, since we have ${\mathcal{D}}\subseteq{\mathcal{C}}_{\mathcal{A}}$ it suffices to show that a cofibrant $X\in{\mathcal{C}}_{\mathcal{A}}$ lies in ${}^{\perp}{\mathcal{W}}_{\mathcal{S}}$ if and only if it lies in ${\mathcal{C}}_{\mathcal{S}}$, in order to obtain that ${\mathcal{D}}={\mathcal{C}}_{\mathcal{A}}\cap{\mathcal{C}}_{\mathcal{S}}$. This is now analogous to (i): we have by 2.16 that $X\in{{}^{\perp}}{\mathcal{W}}_{\mathcal{S}}$ if and only if the image of $X$ in ${\operatorname{Ho}}({\mathcal{A}})$ lies in ${}^{\ddagger}{\mathcal{S}}={\operatorname{Ker}}(\mathbf{L}{\mathsf{L}})$, i.e. if and only if $X\in{\mathcal{C}}_{\mathcal{S}}$, as required. This concludes the proof that the abelian model structure ${\mathcal{A}}^{p}_{\mathcal{S}}$ exists and is cofibrantly generated. The last part is completely dual to (i). ∎
The construction of these model structures now give us the input necessary to apply the techniques of Becker in [7]. We then obtain the following result.

### v4

@@ 4.1. Constructing model structures from a recollement
(4.1) ${{\mathcal{S}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{i}$${{\operatorname{Ho}}({\mathcal{A}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\mathsf{\pi_{\mathcal{S}}}}$$\scriptstyle{{\mathsf{L}}_{{\mathcal{S}}}}$$\scriptstyle{{\mathsf{R}}_{{\mathcal{S}}}}$${{\operatorname{Ho}}({\mathcal{A}})/{\mathcal{S}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\ell_{\mathcal{S}}}$$\scriptstyle{r_{\mathcal{S}}}$
where none of the functors are a priori assumed to be the total derived functor of some Quillen functors.
We recall that the class of trivially fibrant objects in an injective abelian model category equals the class $\mathrm{Inj}({\mathcal{A}})$ of injective objects, and dually the class of trivially cofibrant objects in a projective abelian model category equals the class $\mathrm{Proj}({\mathcal{A}})$ projective objects (cf. [7, Corollary 1.1.9]).
@@
Let ${\mathcal{A}}$ be a Grothendieck abelian category equipped with a cofibrantly generated, abelian model structure which is either injective or projective. Suppose that ${\mathcal{S}}$ is a thick subcategory of ${\operatorname{Ho}}({\mathcal{A}})$ satisfying 4.1. Let ${\mathcal{W}}_{\mathcal{S}}$, resp. ${\mathcal{F}}_{\mathcal{S}}$, resp. ${\mathcal{C}}_{\mathcal{S}}$, denote the class of objects of ${\mathcal{A}}$ whose image in ${\operatorname{Ho}}({\mathcal{A}})$ lies in ${\mathcal{S}}$, resp. ${\operatorname{Ker}}({\mathsf{R}}_{{\mathcal{S}}})$, resp. ${\operatorname{Ker}}({\mathsf{L}}_{{\mathcal{S}}})$.
@@
We briefly explain how our results in this section complement the existing literature on abelian model categories and recollements.
  1. (i)
@@
  2. (ii)
    It came to our attention after this paper first appeared that Gillespie proved quite generally in [12, Theorem A(3)] that recollements in homotopy categories of injective/projective abelian model categories do arise from model structures and Quillen adjunctions. This very nice result was also inspired by the aforementioned techniques of Becker, however he never discusses the cofibrant generation of these model structures and the fibrant/cofibrant objects in his cotorsion pairs are only defined as the orthogonal complements of the trivial objects. 4.2 above may be viewed as a refined version of his result in the setting of 4.1.
@@
Since distinguished triangles arise from short exact sequences, the result follows immediately from the fact that ${\mathcal{S}}$, ${\operatorname{Ker}}({\mathsf{R}}_{{\mathcal{S}}})$, and ${\operatorname{Ker}}({\mathsf{L}}_{{\mathcal{S}}})$ are thick subcategories of the triangulated category ${\operatorname{Ho}}({\mathcal{A}})$. ∎
@@
Suppose that we are in 4.1. When ${\mathcal{A}}$ is injective then ${\mathcal{W}}_{\mathcal{S}}\cap{\mathcal{F}}_{\mathcal{S}}={\mathcal{W}}_{\mathcal{A}}$, so that ${\mathcal{W}}_{\mathcal{S}}\cap{\mathcal{F}}_{\mathcal{A}}\cap{\mathcal{F}}_{\mathcal{S}}={\mathcal{W}}_{\mathcal{A}}\cap{\mathcal{F}}_{\mathcal{A}}=\mathrm{Inj}({\mathcal{A}})$. Dually, when ${\mathcal{A}}$ is projective then ${\mathcal{W}}_{\mathcal{S}}\cap{\mathcal{C}}_{\mathcal{S}}={\mathcal{W}}_{\mathcal{A}}$, so that ${\mathcal{W}}_{\mathcal{S}}\cap{\mathcal{C}}_{\mathcal{A}}\cap{\mathcal{C}}_{\mathcal{S}}={\mathcal{W}}_{\mathcal{A}}\cap{\mathcal{C}}_{\mathcal{A}}=\mathrm{Proj}({\mathcal{A}})$.
@@
We only prove the injective case, the other one being dual. Note that ${\mathcal{W}}_{\mathcal{A}}\cap{\mathcal{F}}_{\mathcal{A}}=\mathrm{Inj}({\mathcal{A}})$ is immediate from the injectivity of the model structure on ${\mathcal{A}}$, so that the first equality immediately implies the second. For the first equality, we have ${\mathcal{W}}_{\mathcal{A}}\subseteq{\mathcal{W}}_{\mathcal{S}}\cap{\mathcal{F}}_{\mathcal{S}}$ by definition of the right hand side. Conversely, assume that $X\in{\mathcal{W}}_{\mathcal{S}}\cap{\mathcal{F}}_{\mathcal{S}}$. Then $X\cong i(Y)$ in ${\operatorname{Ho}}({\mathcal{A}})$, for some $Y\in{\mathcal{S}}$, and $X\in{\operatorname{Ker}}({\mathsf{R}}_{{\mathcal{S}}})$. But then we obtain from 2.4 that $Y=0$ in ${\operatorname{Ho}}({\mathcal{A}})$, and hence that $X\in{\mathcal{W}}_{\mathcal{A}}$ as required. ∎
Let ${\mathcal{A}}$ be a Grothendieck abelian category. Recall that a cotorsion pair $({\mathcal{D}},{\mathcal{E}})$ on ${\mathcal{A}}$ is said to be cogenerated by a set $\mathcal{G}\subseteq{\mathcal{D}}$ if ${\mathcal{E}}=\mathcal{G}^{\perp}$. Furthermore, by definition ${\mathcal{A}}$ has a generating set $\mathscr{S}$ (or equivalently a generator), which is a set of objects of ${\mathcal{A}}$ such that every object of ${\mathcal{A}}$ is a quotient of a direct sum of objects in $\mathscr{S}$. The cotorsion pair $({\mathcal{D}},{\mathcal{E}})$ is called small if it is cogenerated by a set and if ${\mathcal{D}}$ contains a generating set for ${\mathcal{A}}$. We note that small cotorsion pairs are automatically functorially complete and that an abelian model structure $({\mathcal{C}},{\mathcal{F}},{\mathcal{W}})$ on ${\mathcal{A}}$ is cofibrantly generated if and only if both $({\mathcal{C}},{\mathcal{F}}\cap{\mathcal{W}})$ and $({\mathcal{C}}\cap{\mathcal{W}},{\mathcal{F}})$ are small, cf. [17, Corollary 6.6] and [7, Proposition 1.2.7].
We will also need a notion of generating set for a triangulated category ${\mathcal{T}}$. A set of objects $\mathcal{G}$ of ${\mathcal{T}}$ will be called a set of weak generators if an object $X\in{\mathcal{T}}$ is zero if and only if ${\operatorname{Hom}}_{\mathcal{T}}(G,\Sigma^{n}X)=0$ for all $n\in{\mathbb{Z}}$ and all $G\in\mathcal{G}$. The relevance of this notion for us is the result [16, Theorem 7.3.1], which in particular implies that the homotopy category ${\operatorname{Ho}}({\mathcal{A}})$ of a hereditary, abelian, cofibrantly generated model category ${\mathcal{A}}$ has a set of weak generators given by the cokernels of the generating cofibrations. The proof of loc. cit. requires some input from the homotopy theory of simplicial sets, so we deemed it worthwhile to provide a more elementary proof in the case where ${\mathcal{A}}$ is the module category of a Gorenstein ring. In fact we obtain a stronger result than weak generation.
@@
We first note that 4.5 implies that $({\mathcal{A}},{\mathcal{W}}_{\mathcal{S}}\cap{\mathcal{F}}_{\mathcal{A}}\cap{\mathcal{F}}_{\mathcal{S}})=({\mathcal{A}},\mathrm{Inj}({\mathcal{A}}))$ in case (i) and $({\mathcal{W}}_{\mathcal{S}}\cap{\mathcal{C}}_{\mathcal{A}}\cap{\mathcal{C}}_{\mathcal{S}},{\mathcal{A}})=(\mathrm{Proj}({\mathcal{A}}),{\mathcal{A}})$ in case (ii) are small cotorsion pairs, cf. [7, Example 1.2.6]. Since ${\mathcal{W}}_{\mathcal{S}}$ is thick by 4.4, we are only left to show that $({\mathcal{W}}_{\mathcal{S}},{\mathcal{F}}_{\mathcal{A}}\cap{\mathcal{F}}_{\mathcal{S}})$ and $({\mathcal{C}}_{\mathcal{A}}\cap{\mathcal{C}}_{\mathcal{S}},{\mathcal{W}}_{\mathcal{S}})$ are small cotorsion pairs in order to get our claimed model structures (cf. 2.7 and the discussion preceding 4.6).
@@
(i) We need to construct a set $\mathcal{G}$ of objects of ${\mathcal{A}}$ such that ${\mathcal{F}}_{\mathcal{A}}\cap{\mathcal{F}}_{\mathcal{S}}=\mathcal{G}^{\perp}$ and ${\mathcal{W}}_{\mathcal{S}}={{}^{\perp}}(\mathcal{G}^{\perp})$. By assumption on ${\mathcal{A}}$, we have ${\mathcal{F}}_{\mathcal{A}}=\mathcal{G}_{1}^{\perp}$ for some set $\mathcal{G}_{1}$ of objects in ${\mathcal{W}}_{\mathcal{A}}$. By enlarging $\mathcal{G}_{1}$ if necessary, we may assume that it contains a generating set for ${\mathcal{A}}$, since $({\mathcal{W}}_{\mathcal{A}},{\mathcal{F}}_{\mathcal{A}})$ is small. We now let $\mathscr{S}$ be a set of weak generators for ${\operatorname{Ho}}({\mathcal{A}})$ and, for each $S\in\mathscr{S}$ and each $n\in{\mathbb{Z}}$, we choose $X_{n,S}\in{\mathcal{A}}={\mathcal{C}}_{\mathcal{A}}$ such that $X_{n,S}\cong\Sigma^{-n}i{\mathsf{L}}_{{\mathcal{S}}}(S)$ in ${\operatorname{Ho}}({\mathcal{A}})$ and let $\mathcal{G}_{2}=\{X_{n,S}\mid S\in\mathscr{S}\text{ and }n\in{\mathbb{Z}}\}$. Then we claim that $\mathcal{G}=\mathcal{G}_{1}\cup\mathcal{G}_{2}$ satisfies our requirements. To see this, first note that if $Y\in\mathcal{G}^{\perp}$ then in particular $Y\in\mathcal{G}_{1}^{\perp}={\mathcal{F}}_{\mathcal{A}}$. So in order to obtain ${\mathcal{F}}_{\mathcal{A}}\cap{\mathcal{F}}_{\mathcal{S}}=\mathcal{G}^{\perp}$ it suffices to show that for $Y\in{\mathcal{F}}_{\mathcal{A}}$, we have $Y\in\mathcal{G}_{2}^{\perp}$ if and only if $Y\in{\mathcal{F}}_{\mathcal{S}}$. Indeed, by 2.16 we have
@@
$Y\in\mathcal{G}_{2}^{\perp}$ $\iff{\operatorname{Hom}}_{{\operatorname{Ho}}({\mathcal{A}})}(\Sigma^{-n}i{\mathsf{L}}_{{\mathcal{S}}}(S),\Sigma Y)=0\text{ for all }n\in{\mathbb{Z}}\text{ and }S\in\mathscr{S}$
$\iff{\operatorname{Hom}}_{{\operatorname{Ho}}({\mathcal{A}})}(S,\Sigma^{n+1}i{\mathsf{R}}_{{\mathcal{S}}}(Y))=0\text{ for all }n\in{\mathbb{Z}}\text{ and }S\in\mathscr{S}$
$\iff i{\mathsf{R}}_{{\mathcal{S}}}(Y)=0\iff Y\in{\mathcal{F}}_{\mathcal{S}}$
@@
as required. Next, note that every object of ${\mathcal{F}}_{\mathcal{S}}$ is isomorphic in ${\operatorname{Ho}}({\mathcal{A}})$ to an object of ${\mathcal{F}}_{\mathcal{A}}\cap{\mathcal{F}}_{\mathcal{S}}$. Thus we see by applying 2.16 again that an object $X\in{\mathcal{A}}$ lies in ${}^{\perp}({\mathcal{F}}_{\mathcal{A}}\cap{\mathcal{F}}_{\mathcal{S}})$ if and only if the image of $X$ in ${\operatorname{Ho}}({\mathcal{A}})$ lies in ${}^{\ddagger}{\operatorname{Ker}}({\mathsf{R}}_{{\mathcal{S}}})={\mathcal{S}}$.
@@
This shows that $({\mathcal{W}}_{\mathcal{S}},{\mathcal{F}}_{\mathcal{A}}\cap{\mathcal{F}}_{\mathcal{S}})$ is a small cotorsion pair, and thus we have that ${\mathcal{A}}^{i}_{\mathcal{S}}$ defines an injective, cofibrantly generated, abelian model structure as required. By 4.5 we have that the trivial fibrations in ${\mathcal{A}}$ and in ${\mathcal{A}}^{i}_{\mathcal{S}}$ are the same, and thus ${\operatorname{id}}:{\mathcal{A}}^{i}_{\mathcal{S}}\to{\mathcal{A}}$ is right Quillen. Furthermore, since the trivially fibrant objects in both model structures are the same we have that $\mathbf{R}{\operatorname{id}}:{\operatorname{Ho}}({\mathcal{A}}^{i}_{\mathcal{S}})\to{\operatorname{Ho}}({\mathcal{A}})$ identifies, after restricting to the cofibrant-fibrant objects, with the canonical embedding $({\mathcal{F}}_{\mathcal{S}}\cap{\mathcal{F}}_{\mathcal{A}})/\mathrm{Inj}({\mathcal{A}})\hookrightarrow{\mathcal{F}}_{\mathcal{A}}/\mathrm{Inj}({\mathcal{A}})$ with image ${\operatorname{Ker}}({\mathsf{R}}_{{\mathcal{S}}})\simeq{\operatorname{Ho}}({\mathcal{A}})/{\mathcal{S}}$, where the latter equivalence is induced by $r_{\mathcal{S}}$. Thus, after these identifications, the left adjoint of $\mathbf{R}{\operatorname{id}}$ is naturally isomorphic to $\pi_{\mathcal{S}}$.
@@
(ii) As above, we let $\mathscr{S}$ be a set of weak generators for ${\operatorname{Ho}}({\mathcal{A}})$. Then for each $S\in\mathscr{S}$ and each $n\in{\mathbb{Z}}$, we choose $X_{n,S}\in{\mathcal{C}}_{\mathcal{A}}$ such that $X_{n,S}\cong\Sigma^{-n}\ell_{\mathcal{S}}\pi_{\mathcal{S}}(S)$ in ${\operatorname{Ho}}({\mathcal{A}})$ and let $\mathcal{G}=\{X_{n,S}\mid S\in\mathscr{S}\text{ and }n\in{\mathbb{Z}}\}$. Since the model structure on ${\mathcal{A}}$ is projective, there is a projective generator $Z$ for ${\mathcal{A}}$ and we may apply [17, Corollary 6.9] to see that $\mathcal{G}$ cogenerates a small cotorsion pair $({\mathcal{D}},{\mathcal{E}})$, where ${\mathcal{D}}$ is the closure of $\mathcal{G}\cup\{Z\}$ under direct summands and certain colimits called transfinite compositions. As ${\mathcal{C}}_{\mathcal{A}}$ is closed under those (cf. Lemma 6.2 in loc. cit.), it follows that ${\mathcal{D}}\subseteq{\mathcal{C}}_{\mathcal{A}}$. We first show that ${\mathcal{E}}=\mathcal{G}^{\perp}={\mathcal{W}}_{\mathcal{S}}$. This is similar to the argument in (i). As every object of $\mathcal{G}$ is cofibrant, we may apply 2.16 to see that
@@
as required, where the last equivalence holds because ${\operatorname{Ker}}(\pi_{\mathcal{S}})={\mathcal{S}}$. Next, since we have ${\mathcal{D}}\subseteq{\mathcal{C}}_{\mathcal{A}}$ it suffices to show that a cofibrant $X\in{\mathcal{C}}_{\mathcal{A}}$ lies in ${}^{\perp}{\mathcal{W}}_{\mathcal{S}}$ if and only if it lies in ${\mathcal{C}}_{\mathcal{S}}$, in order to obtain that ${\mathcal{D}}={\mathcal{C}}_{\mathcal{A}}\cap{\mathcal{C}}_{\mathcal{S}}$. This is now analogous to (i): we have by 2.16 that $X\in{{}^{\perp}}{\mathcal{W}}_{\mathcal{S}}$ if and only if the image of $X$ in ${\operatorname{Ho}}({\mathcal{A}})$ lies in ${}^{\ddagger}{\mathcal{S}}={\operatorname{Ker}}({\mathsf{L}}_{{\mathcal{S}}})$, i.e. if and only if $X\in{\mathcal{C}}_{\mathcal{S}}$, as required. This concludes the proof that the abelian model structure ${\mathcal{A}}^{p}_{\mathcal{S}}$ exists and is cofibrantly generated. The last part is completely dual to (i). ∎
The construction of these model structures now gives us the input necessary to apply the techniques of Becker in [7]. We then obtain the following result.